OpenStudy (anonymous):

**help me please** -Simplify the following equation:

OpenStudy (anonymous):

what is the equation

OpenStudy (anonymous):

\[(y-2)(3y ^{2}+y-5)\]

OpenStudy (anonymous):

alright to simplify it you have to distribute

OpenStudy (zarkon):

"Simplify the following equation:" that is not an equation

OpenStudy (jhannybean):

Use the distributive property for both values \(y\) and \(-2\)

OpenStudy (jhannybean):

\[(y-2)(3y ^{2}+y-5)\]\[= (y)(3y^2) +(y)(y) +(y)(-5) +(-2)(3y^2)+(-2)(y) +(-2)(-5)\]

OpenStudy (anonymous):

anddruy

OpenStudy (jhannybean):

simplify that, combine like terms (if there are any) and you'll find your answer :)

OpenStudy (anonymous):

@Jhannybean \[-2y ^{2}-7y+10\]

OpenStudy (anonymous):

is that correct ?

OpenStudy (jhannybean):

No.

OpenStudy (anonymous):

um no

OpenStudy (anonymous):

lol seriously! ughh Im so bad at math! ok let me try again. thank yall for helping

OpenStudy (anonymous):

is the first part \[x^{3} or x ^{2}\]

OpenStudy (jhannybean):

when you multiply two of the same variables with eachother that have different powers, the powers add together. ex: \(x^4 \cdot x^7 = x^{4+7} = x^{11}\)

OpenStudy (anonymous):

ok is the answer \[3y ^{3}-5y ^{2} -7y+10\] or am I close ?

OpenStudy (jhannybean):

\[3y^2 +(y^2 -6y^2) +(-5y -2y)+10\]\[=3y^2 -5y^2 -7y +10 ~ \checkmark\] Good job.

OpenStudy (jhannybean):

Remember to medal the user that helped you the most :)

OpenStudy (anonymous):

@Jhannybean you got it ! can you help me with another problem I like that you didn't just give me the answer but helped.

OpenStudy (jhannybean):

I can try